OpenStudy (anonymous):

Integral of x^2/(x^2+2)

OpenStudy (anonymous):

\[\int\limits X^2dx/(X^2+2)\]

OpenStudy (anonymous):

\[\int\limits \frac{ x^2}{ X^2+2 }\]

OpenStudy (perl):

first do long division

OpenStudy (anonymous):

how can you tell that?

OpenStudy (anonymous):

I will, one sec

OpenStudy (anonymous):

because of similarity?

OpenStudy (anonymous):

\[\int\limits 1-\frac{ 2 }{ x^2+1 }\]

OpenStudy (perl):

because the exponent of the numerator is the same (or larger than) the denominator, we should do long division

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

from here I didn't get the same answer as written though...

OpenStudy (anonymous):

I got x-arctanX

OpenStudy (perl):

$$ \int\limits 1-\frac{ 2 }{ x^2+\color{red}2 } $$

OpenStudy (anonymous):

yeah. and from here I divide it into two integrals

OpenStudy (perl):

$$ \Large{ \int\limits \frac{ x^2}{ x^2+2 } =\int\limits 1-\frac{ 2 }{ x^2+2 }\\ =x - 2\cdot \int \frac{1}{x^2+2}\\ =x - 2\cdot \int \frac{1}{2(\frac{x^2}{2}+1)}\\ =x - 2\cdot \int \frac{1}{2~( (\frac{x}{\sqrt{2} })^2+1)}\\ =x - \frac{2}{2}\cdot \int \frac{1}{ (\frac{x}{\sqrt{2} })^2+1)}\\ =x - \int \frac{1}{ (\frac{x}{\sqrt{2} })^2+1)}\\ \\~~\\ u = \frac{x}{\sqrt{2}}\\ du = ... } $$ Now apply u substitution

OpenStudy (anonymous):

you're trully brilliant !!! I've got it! Thanks allot helping me out!!!!!!!!

OpenStudy (anonymous):

for*

OpenStudy (anonymous):

Directly substituting a trig expression works nicely as well. Let \(x=\sqrt2\tan t\), then \(dx=\sqrt2\sec^2t\,dt\), then \[\int\frac{x^2}{x^2+2}\,dx=\int\frac{\sqrt2\sec^2t}{2\tan^2t+2}\,dt=\frac{1}{\sqrt2}\int\,dt=\cdots\]